OpenStudy (anonymous):

A Norman window has the shape of a rectangle surmounted by a semicircle. If the perimeter of the window is 31 feet, express the area A of the window as a function of the width x (across the base) of the window.

Directrix (directrix):

Let the height of the rectangle be h and the radius of the semicircle be r. Area of rectangle = (2r)*h = 2rh Area of semicircle = 1/2 pi r^2 A = 2rh + 1/2 pi r^2 but there are two variables, so we'll first try to eliminate one of them. Note that perimeter = 2h + 2r + pi r = 31 so h = 31 - r - pi/2 r = 31 - (1 + pi/2)r Substitute that into our equation for area to get: A = 2r(31-(1+pi/2)r) + 1/2 pi r^2 = 62r - 2r^2 - 1/2 pi r^2 @Yondaime Study this and see what you think. Solution written by @AwmsA

Directrix (directrix):

Question?